OpenStudy (anonymous):

Simplify: (x-3)/(x^2-x-6)

OpenStudy (anonymous):

did you factorise the denominator?

OpenStudy (anonymous):

Try to Factorize x^2-x-6)

OpenStudy (anonymous):

x^2-x-6 = (x-3) (x+2)

OpenStudy (anonymous):

\[\frac{ (x-3 )}{ (x-3) (x+2) }\]

OpenStudy (anonymous):

=\[\frac{ 1 }{ x+2 }\]

OpenStudy (anonymous):

as Yahoo stated, it can be factorised as such, though you'll need to know the buttefly method to factorise it.

OpenStudy (anonymous):